Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-071_2000-077

Mission Operations Days: 2000/071 to 2000/077
March 11, 2000 00:00:00 GMT to March 17, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Band 14H Gain Change (10 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Friday, March 10, 2000

2000/070 20:11:00 – 20:22:49 Real-time Ecal Sweep; Vcal starts 48,43,53,60 8 Vcal steps of 1 DN, Nstep of 22,22,22,10
2000/070 21:39:23 Real-time - Set new gain values for bands 5, 14H, and 16
2000/070 21:59:17 Real-time - Set Blackbody to 290K
2000/070 22:14:29 ATC – OA15: SD/SDSM Open Double

Saturday, March 11, 2000

2000/071 03:11:08 ATC – OA16: SD/SDSM Screened Double
2000/071 08:07:46 ATC – OA16: SD/SDSM Screened Double
2000/071 13:04:25 ATC – OA15: SD/SDSM Open Double
2000/071 16:35:43 – 16:42:11 ATC – OA14:Sector Rotation for entire cavity (DCR Off) Deltas of 0, 1776, 4711, 5643, & 6842
2000/071 18:00:04 ATC – OA16: SD/SDSM Screened Double

Sunday, March 12, 2000

2000/072 03:54:06 ATC – OA16: SD/SDSM Screened Double
2000/072 08:50:44 ATC – OA15: SD/SDSM Open Double
2000/072 14:29:58 – 14:41:48 Real-time Ecal Sweep; Vcal starts 48,43,53,60 8 Vcal steps of 1 DN, Nstep of 20,20,20,8
2000/072 15:26:14 ATC – OA16: SD/SDSM Screened Double
2000/072 15:39:45 – 15:52:23 Real-time Ecal Sweep; Vcal starts 48,43,53,60 8 Vcal steps of 1 DN, Nstep of 18,18,18,6
2000/072 16:23:01 – 16:34:47 Real-time Ecal Sweep; Vcal starts 48,43,53,60 8 Vcal steps of 1 DN, Nstep of 22,22,22,10
2000/072 20:23:52 ATC – OA16: SD/SDSM Screened Double

Monday, March 13, 2000

2000/073 01:19:29 ATC – OA16: SD/SDSM Screened Double
2000/073 02:14:40 ATC – OA19: SRCA Full Radiometric
2000/073 04:37:14 ATC – OA15: SD/SDSM Open Double
2000/073 07:54:59 ATC – OA15: SD/SDSM Screened Double
2000/073 11:12:44 ATC – OA15: SD/SDSM Screened Double
2000/073 13:54:52 – 15:33:45 ATC – OA24: SRCA 10W Radiometric (Rad Feedback Mode)
2000/073 20:30:23 – 22:09:15 ATC – OA24: SRCA 10W Radiometric (Constant Current Mde)

Tuesday, March 14, 2000

2000/074 00:23:45 ATC – OA15: SD/SDSM Open Double
2000/074 03:05:53 – 04:44:45 ATC – OA25: SRCA 1W Radiometric
2000/074 05:20:22 ATC – OA16: SD/SDSM Screened Double
2000/074 16:52:30 ATC – OA16: SD/SDSM Screened Double
2000/074 18:31:22 ATC – OA15: SD/SDSM Open Double
2000/074 20:38:52 Drag Make-up Maneuver 22.777 seconds long
2000/074 22:17:10 Real-time: Reset Band 5 & 16 gains to incorrect (launch) values
2000/074 23:01:31 Real-time: Memory Dump of Table 4, PV Gains

Wednesday, March 15, 2000

2000/075 00:59 – 01:10:42 Real-time: Several Itwk sweeps; Vdet =226 Itwk = 70 through 250 by steps of 10
2000/075 01:37:16 – 01:48:34 Real-time: Several Itwk sweeps; Vdet =225 Itwk = 70 through 250 by steps of 10
2000/075 02:36:34 – 02:48:07 Real-time: Several Itwk sweeps; Vdet =224 Itwk = 70 through 250 by steps of 10
2000/075 02:45:46 ATC – OA16: SD/SDSM Screened Double
2000/075 03:23:05 – 03:34:35 Real-time: Several Itwk sweeps; Vdet =223 Itwk = 70 through 250 by steps of 10
2000/075 04:14:08 – 04:25:50 Real-time: Several Itwk sweeps; Vdet =222 Itwk = 70 through 250 by steps of 10
2000/075 05:28:40 – 05:40:11 Real-time: Several Itwk sweeps; Vdet =221 Itwk = 70 through 250 by steps of 10
2000/075 06:19:35 – 06:31:12 Real-time: Several Itwk sweeps; Vdet =220 Itwk = 70 through 250 by steps of 10
2000/075 07:10:51 – 07:22:25 Real-time: Several Itwk sweeps; Vdet =219 Itwk = 70 through 250 by steps of 10
2000/075 07:42:25 ATC – OA16: SD/SDSM Screened Double
2000/075 08:47:54 – 08:59:31 Real-time: Several Itwk sweeps; Vdet =218 Itwk = 70 through 250 by steps of 10
2000/075 12:42:52 ATC – OA15: SD/SDSM Open Double
2000/075 15:05:19 – 16:02:50 ATC – OA23: SRCA Full Spatial
2000/075 19:14:36 ATC – OA16: SD/SDSM Screened Double

Thursday, March 16, 2000

2000/076 07:00 ATC – Blackbody to 280K
2000/076 12:00 ATC – Blackbody to 285K
2000/076 15:00 ATC – Blackbody to 290K
2000/076 15:00 Real-time – Set Blackbody duty cycle to 100%
2000/076 17:00 ATC – Blackbody to 295K
2000/076 18:45:43 Yaw Maneuver +15 deg (33 min, 54 sec)
2000/076 19:00 ATC – Blackbody to 300K
2000/076 20:24:35 Yaw Maneuver -15 deg (33 min, 54 sec)
2000/076 20:49:41 Real-time – Blackbody to 33%
2000/076 21:00 ATC – Blackbody to 270K
2000/076 21:22:31 Real-time – Reset gain changes for bands 5 & 16 to correct launch values
2000/076 21:36:46 Real-time – Ran MOD_MEM_HALT to clear memory transfer
2000/076 22:10:18 Real-time – Dumped PV Gain Table

Friday, March 17, 2000

2000/077 15:04:20 Real-time Set Blackbody to 290K

This Week's Scheduled MODIS Activities Not Completed:

2000/076 21:27:04 Real-time – Failure during 1st dump attempt of gain table due to ground system error.

Upcoming MODIS Events:

*SRCA activities are scheduled, but inhibited (so they will not execute) until further notice.

Friday, March 17, 2000

2000/077 20:41:03 ATC – OA16: SD/SDSM Screened Double

Saturday, March 18, 2000

2000/078 03:16:35 ATC – OA16: SD/SDSM Open Double
2000/078 09:52:06 ATC – OA16: SD/SDSM Screened Double
2000/078 16:27:38 ATC – OA16: SD/SDSM Screened Double
2000/078 23:03:09 ATC – OA16: SD/SDSM Open Double

Sunday, March 19, 2000

2000/079 05:38:41 ATC – OA16: SD/SDSM Screened Double
2000/079 12:14:12 ATC – OA16: SD/SDSM Screened Double
2000/079 18:49:44 ATC – OA16: SD/SDSM Open Double

Monday, March 20, 2000

2000/080 01:25:15 ATC – OA16: SD/SDSM Screened Double
2000/080 08:00:47 ATC – OA16: SD/SDSM Screened Double
2000/080 14:36:18 ATC – OA16: SD/SDSM Open Double
2000/080 21:11:50 ATC – OA16: SD/SDSM Screened Double

Tuesday, March 21, 2000

2000/081 03:47:21 ATC – OA16: SD/SDSM Screened Double
2000/081 10:22:53 ATC – OA16: SD/SDSM Open Double
2000/081 15:00?? Real-time Set Blackbody duty cycle to 100%
2000/081 15:00 ATC – Blackbody to 315K
2000/081 16:58:24 ATC – OA16: SD/SDSM Screened Double
2000/081 22:00:00 ATC – Blackbody to 270K
2000/081 22:15:58 Real-time Set Blackbody duty cycle to 33%
2000/081 23:33:56 ATC – OA16: SD/SDSM Screened Double

Wednesday, March 22, 2000

2000/082 12:01:13 ATC – OA-23 SRCA Full Spatial
2000/082 19:00:00 ATC – Blackbody to 290K
2000/082 ?? Real-time Ecal sweeps

Thursday, March 23, 2000

2000/083 15:00?? Real-time Set Blackbody duty cycle to 100%
2000/083 15:00:00 ATC – Blackbody to 315K
2000/083 22:00:00 ATC – Blackbody to 270K
2000/083 22:00?? Real-time Set Blackbody duty cycle to 33%

Friday, March 24, 2000

2000/084 08:28:01 ATC – OA-22 SRCA Full Spectral Part 1
2000/084 10:22:46 ATC – OA-22 SRCA Full Spectral Part 2
2000/084 11:45:11 ATC – OA-22 SRCA Full Spectral Part 3
2000/084 13:00 ATC – Blackbody to 290K
2000/084 13:40:31 ATC – OA-22 SRCA Full Spectral Part 4
2000/084 16:52:51 ATC – 1st Roll Manuever to view moon in space-view port (-19.5 deg)
Sector rotation to –3072, DCR off, Formatter Day rate
2000/084 19:00?? Real-time Set SMIR Itwk / Vdet to ?? / ??
2000/084 20:09:57 ATC – 2nd Roll Manuever to view moon in space-view port (-19.0 deg)
Sector rotation to –3072, DCR off, Formatter Day rate
2000/084 21:00?? Real-time Set SMIR Itwk / Vdet to ?? / ??

Fri, March 24, 2000 - MODIS Roll Maneuver (-19 deg)
Tue, April 25, 2000 - MODIS Yaw Maneuvers
Wed, April 26, 2000 - MODIS Yaw Maneuvers
Thur, April 27, 2000 - Inclination Maneuver
Fri, April 28, 2000 - Post-inclination Drag Makeup Maneuver

MODIS Anomalies:

None.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open.

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

The transition to normal operations continues as the FOT has begun running our morning status meeting as Project Leads begin the handover.

Limited Life Item Status:

SRCA 10W Lamp #1: 146.0 of 500 hours
SRCA 10W Lamp #2: 124.5 of 500 hours
SRCA 10W Lamp #3: 133.7 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 545.5 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1121 of 3022 Movements
Nadir Aperture Door: 530 of 1316 Movements
Space View Door: 436 of 1316 Movements